Number of moles of HCI (MM=36.46 g/mol)needed to completely react with Ca(OH)₂.The molar mass of Ca(OH)₂ is 74.10 g/mol.


1
Expert's answer
2021-06-21T07:28:49-0400

Ca(OH)2+2HCl=CaCl2+2H2O

v(Ca(OH)2)=v/2(HCl)=0.5mol

v(HCl)=1mol